If peter pays $84 to buy a watch after a discount of 20% find the original price of the watch
1
Expert's answer
2019-12-12T10:29:33-0500

Let $x be the original price of the watch. Then $(x-84) is the sum of a discount.

So we obtain the equation:

"x - 84 = 0.2 x; \newline 
 x - 0.2x = 84; \newline
0.8x = 84; \newline
x =\frac{84}{0.8}; \newline
x = 105"

Thus, the original price of the watch is $105.